# Which Optimizer Works Best for Physics-Informed Neural Networks and Kolmogorov-Arnold Networks?

### v4

@@ 3 Optimizing Data-Driven DeepONet with Self-Scaled Optimizers
Figure 29: Double-Precision DeepONet Comparison of predicted and true physical fields using the SSBroyden optimizer: (a) displacement $u$, and (b) damage $\alpha$.
In this section, we extend our study to evaluate the performance of second-order optimizers BFGS, SSBFGS, and SSBroyden for DeepONet in a purely data-driven setting. Deep Operator Networks (DeepONets) are neural architectures designed to learn operators—mappings between function spaces—and are particularly effective for solving parametric PDEs [94]. The model consists of two subnetworks, a branch network that encodes input functions sampled at discrete sensor points, and a trunk network that encodes the spatial coordinates at which the output is evaluated. The final output is obtained via an inner product between the branch and trunk outputs, approximating the operator mapping $\mathcal{G}:F\mapsto G$. For full architectural and theoretical details, see [94].
@@
Figure 30: Double-Precision DeepONet on the 1D dataset. (a) Loss evolution as well as relative $l_{2}$ error over optimizer iterations. (b) Relative $l_{2}$ error in damage variable $\alpha$. (c) Relative $l_{2}$ error in displacement field $u$. The table summarizes training performance for different optimization strategies.
@@
Table 17: Double-Precision DeepONet on the 1D dataset: Comparison of optimizer configurations for the 1D fracture problem. Relative $l_{2}$ errors are reported for displacement $u$ and damage field $\alpha$, and total training time is shown in seconds.
Figure 29 compares (a) the predicted damage field $\alpha$ with the true $\alpha$, and (b) the predicted displacement field $u$ with the true $u$, for 5 testing displacement steps. Figure 30 presents (a) the training loss for four optimizers—Adam, BFGS, SSBFGS, and SSBroyden—where the loss function includes contributions from both the damage field $\alpha$ and the displacement field $u$. It also shows the relative error during training over iterations, demonstrating that SSBroyden and SSBFGS achieve a relative error of $10^{-3}$ with fewer iterations compared to Adam and BFGS. Subfigures (b) and (c) show the absolute error in predicting the damage field $\alpha$ and displacement field $u$, respectively, over the spatial domain. Notably, the error from SSBroyden and SSBFGS remains below 0.0025 for $\alpha$ and around 0.005 for $u$, especially near $x=1$, where the phase field transitions sharply from 0 to 1. In contrast, other optimizers exhibit increased errors at this critical region, highlighting the superior stability and accuracy of SSBroyden. Table 8 summarizes the relative errors and training times for several optimization strategies applied to the same network architecture. In all cases, the network size and data remain fixed. Case 1 uses Adam for 15000 iterations. Cases 2 and 3 combine a warm-up phase using Adam with 6,000 iterations of second-order optimizers. Specifically, Case 2 uses only 10 iterations of Adam before switching to BFGS, SS-BFGS, or SSBroyden, while Case 3 uses a longer warm-up of 100 iterations. The results show that second-order methods, particularly SSBroyden, consistently yield lower errors. While SSBroyden requires slightly longer training times, it achieves the best accuracy across both Case 2 and Case 3 configurations.
@@ 4 Summary
In this study, we conducted a comprehensive comparison of multiple second-order optimizers for PINNs and PIKANs, focusing on BFGS, SSBFGS, SSBroyden, and L-BFGS with Wolfe line-search conditions, as well as BFGS with Backtracking line search and trust-region methods. Utilizing the optax and optimistix library in JAX, we systematically evaluated these optimization techniques across a range of PDEs. We note that we obtained state-of-the-art results by only employing the best optimizers without fine-tuning our loss functions with self-adaptive or attention-based weights or any other enhancements. We simply pursued a straightforward application of good optimization solvers in order to demonstrate that the big bottleneck for PINNs or PIKANs is the optimization error.
@@
Our initial investigation focused on the Burgers equation, where we evaluated the impact of combining first-order optimizers such as Adam with second-order optimizers such as BFGS and SSBroyden. We then extended our analysis to more challenging PDEs, including the Allen-Cahn equation, the Kuramoto-Sivashinsky equation, and the Ginzburg-Landau equation, providing a detailed evaluation of efficiency and accuracy. The results demonstrate that advanced quasi-Newton methods, particularly SSBroyden, SSBFGS and BFGS with Wolfe line-search, significantly improve the convergence rate and accuracy of PINNs, especially for complex and stiff PDEs. Across all cases, SSBroyden consistently outperformed SSBFGS and BFGS, achieving faster convergence and exhibiting greater robustness in handling complex optimization landscapes. These findings highlight the effectiveness of quasi-Newton methods in accelerating the training of PINNs and enhancing numerical stability. Furthermore, we extended our analysis to PIKANs, replacing traditional multilayer perceptron architectures with KANs utilizing Chebyshev polynomials. In this setting, SSBroyden continued to demonstrate superior optimization performance, reinforcing its robustness across diverse network architectures. Additionally, our study emphasized the importance of implementing PINNs and PIKANs with double-precision arithmetic, which improves numerical stability and enhances optimization efficiency across all tested scenarios. Moreover, we evaluated the performance of second-order optimizers —BFGS, SSBFGS, and SSBroyden— for DeepONet in a purely data-driven setting. Specifically, we studied crack nucleation in a one-dimensional homogeneous bar subjected to increasing prescribed displacements.
@@
In Appendix B, we provide a detailed comparison of the number of iterations required to minimize the Rosenbrock function using various optimizers. This analysis highlights how increasing the dimensionality affects the convergence behavior of second-order methods such as BFGS and SSBroyden. We observed that these methods remain efficient across dimensions, while first-order optimizers like ADAM require significantly more iterations and fail to achieve comparable accuracy. The errors reported in this study show that using SSBroyden with the Wolfe line search leads to significantly lower relative $l_{2}$ errors in all benchmark problems tested compared to the state-of-the-art SOAP method [56]. For instance, in the Burgers equation, our approach reduces the error from $8.06\times 10^{-6}$ to $4.19\times 10^{-8}$. In the Kuramoto–Sivashinsky equation, we achieve an error of $2.65\times 10^{-5}$ over the full time interval $t\in[0,1]$, whereas SOAP reports $3.86\times 10^{-2}$ over a shorter interval $t\in[0,0.8]$. Similar improvements are observed for the Allen–Cahn and Ginzburg–Landau equations (see Table 18). A head-to-head comparison of SSBroyden shows that, unlike other second-order optimizers, SSBroyden achieves comparable accuracy with dramatically smaller networks—using just five hidden layers of 30 neurons for the Kuramoto-Sivashinsky equation and three hidden layers of 30 neurons for the Allen–Cahn equation—resulting in substantially reduced training time.
@@ Acknowledgments
Appendix A Lorenz system
@@
Appendix B Multi-dimensional Rosenbrock function

### v5

@@ 3 Optimizing Data-Driven DeepONet with Self-Scaled Optimizers
Figure 24: Double-Precision DeepONet. Comparison of predicted and true physical fields using the SSBroyden optimizer: (a) displacement $u$, and (b) damage $\alpha$.
In this section, we extend our study to evaluate the performance of quasi-Newton optimizers BFGS, SSBFGS, and SSBroyden for DeepONet in a purely data-driven setting. Deep Operator Networks (DeepONets) are neural architectures designed to learn operators—mappings between function spaces—and are particularly effective for solving parametric PDEs [75]. The model consists of two subnetworks, a branch network that encodes input functions sampled at discrete sensor points, and a trunk network that encodes the spatial coordinates at which the output is evaluated. The final output is obtained via an inner product between the branch and trunk outputs, approximating the operator mapping $\mathcal{G}:F\mapsto G$. For full architectural and theoretical details, see [75].
@@
Figure 25: Double-Precision DeepONet on one-dimensional homogeneous bar. (a) Loss evolution as well as relative $l_{2}$ error over optimizer iterations. (b) Relative $l_{2}$ error in damage variable $\alpha$. (c) Relative $l_{2}$ error in displacement field $u$. The table summarizes training performance for different optimization strategies.
@@
Table 15: Double-Precision DeepONet on one-dimensional homogeneous bar.: Comparison of optimizer configurations for the 1D fracture problem. Relative $l_{2}$ errors are reported for displacement $u$ and damage field $\alpha$, and total training time is shown in seconds.
Figure 24 compares (a) the predicted damage field $\alpha$ with the true $\alpha$, and (b) the predicted displacement field $u$ with the true $u$, for 5 testing displacement steps. Figure 25 presents (a) the training loss for four optimizers—Adam, BFGS, SSBFGS, and SSBroyden—where the loss function includes contributions from both the damage field $\alpha$ and the displacement field $u$. It also shows the relative error during training over iterations, demonstrating that SSBroyden and SSBFGS achieve a relative error of $10^{-3}$ with fewer iterations compared to Adam and BFGS. Subfigures (b) and (c) show the absolute error in predicting the damage field $\alpha$ and displacement field $u$, respectively, over the spatial domain. Notably, the error from SSBroyden and SSBFGS remains below 0.0025 for $\alpha$ and around 0.005 for $u$, especially near $x=1$, where the phase field transitions sharply from 0 to 1. In contrast, other optimizers exhibit increased errors at this critical region, highlighting the superior stability and accuracy of SSBroyden. Table 6 summarizes the relative errors and training times for several optimization strategies applied to the same network architecture. In all cases, the network size and data remain fixed. Case 1 uses Adam for 15000 iterations. Cases 2 and 3 combine a warm-up phase using Adam with 6,000 iterations of quasi-Newton optimizers. Specifically, Case 2 uses only 10 iterations of Adam before switching to BFGS, SS-BFGS, or SSBroyden, while Case 3 uses a longer warm-up of 100 iterations. The results show that quasi-Newton methods, particularly SSBroyden, consistently yield lower errors. While SSBroyden requires slightly longer training times, it achieves the best accuracy across both Case 2 and Case 3 configurations.
@@ 4 Summary
In this study, we conducted a comprehensive comparison of multiple quasi-Newton optimizers for PINNs and PIKANs, focusing on BFGS, SSBFGS, SSBroyden, and L-BFGS with Wolfe line-search conditions, as well as BFGS with Backtracking line search and trust-region methods. Utilizing the optax and optimistix library in JAX, we systematically evaluated these optimization techniques across a range of PDEs. We note that we obtained state-of-the-art results by only employing the best optimizers without fine-tuning our loss functions with self-adaptive or attention-based weights or any other enhancements. We simply pursued a straightforward application of good optimization solvers in order to demonstrate that the big bottleneck for PINNs or PIKANs is the optimization error.
@@
Our initial investigation focused on the Burgers equation, where we evaluated the impact of combining first-order optimizers such as Adam with quasi-Newton optimizers such as BFGS and SSBroyden. We then extended our analysis to more challenging PDEs, including the Allen-Cahn equation, the Kuramoto-Sivashinsky equation, and the Ginzburg-Landau equation, providing a detailed evaluation of efficiency and accuracy. The results demonstrate that advanced quasi-Newton methods, particularly SSBroyden, SSBFGS and BFGS with Wolfe line-search, significantly improve the convergence rate and accuracy of PINNs, especially for complex and stiff PDEs. Across all cases, SSBroyden consistently outperformed SSBFGS and BFGS, achieving faster convergence and exhibiting greater robustness in handling complex optimization landscapes. These findings highlight the effectiveness of quasi-Newton methods in accelerating the training of PINNs and enhancing numerical stability. Furthermore, we extended our analysis to PIKANs, replacing traditional multilayer perceptron architectures with KANs utilizing Chebyshev polynomials. In this setting, SSBroyden continued to demonstrate superior optimization performance, reinforcing its robustness across diverse network architectures. Additionally, our study emphasized the importance of implementing PINNs and PIKANs with double-precision arithmetic, which improves numerical stability and enhances optimization efficiency across all tested scenarios. Moreover, we evaluated the performance of quasi-Newton optimizers —BFGS, SSBFGS, and SSBroyden— for DeepONet in a purely data-driven setting. Specifically, we studied crack nucleation in a one-dimensional homogeneous bar subjected to increasing prescribed displacements.
@@
In Appendix D, we provide a detailed comparison of the number of iterations required to minimize the Rosenbrock function using various optimizers. This analysis highlights how increasing the dimensionality affects the convergence behavior of quasi-Newton methods such as BFGS and SSBroyden. We observed that these methods remain efficient across dimensions, while first-order optimizers like ADAM require significantly more iterations and fail to achieve comparable accuracy. The errors reported in this study show that using SSBroyden with the Wolfe line search leads to significantly lower relative $l_{2}$ errors in all benchmark problems tested compared to the state-of-the-art SOAP method [56]. For instance, in the Burgers equation, our approach reduces the error from $8.06\times 10^{-6}$ to $4.19\times 10^{-8}$. In the Kuramoto–Sivashinsky equation, we achieve an error of $2.65\times 10^{-5}$ over the full time interval $t\in[0,1]$, whereas SOAP reports $3.86\times 10^{-2}$ over a shorter interval $t\in[0,0.8]$. Similar improvements are observed for the Allen–Cahn and Ginzburg–Landau equations (see Table 16). A head-to-head comparison of SSBroyden shows that, unlike other quasi-Newton optimizers, SSBroyden achieves comparable accuracy with dramatically smaller networks—using just five hidden layers of 30 neurons for the Kuramoto-Sivashinsky equation and three hidden layers of 30 neurons for the Allen–Cahn equation—resulting in substantially reduced training time.
@@ Acknowledgments
Appendix A Search Methods and Trust-Region Techniques
In this section, we provide a detailed overview of search methods and trust-region techniques in the context of quasi-Newton algorithms, focusing on their role in achieving efficient and reliable convergence.
A.0.1 Line-search methods
$\min_{\alpha>0}f(\mathbf{x}_{k}+\alpha\mathbf{p}_{k}).$ (17)
$\mathbf{x}_{k+1}=\mathbf{x}_{k}+\alpha_{k}\mathbf{p}_{k}.$ (18)
Wolfe conditions
which commonly receive the names of Armijo and sufficient decrease conditions, respectively. The quantities $c_{1}$ and $c_{2}$ are constants that should follow $0<c_{1}<c_{2}<1$. While the first condition ensures a sufficient decrease in the function $f$, the second one rules out unacceptably short steps. The latter inequality is commonly replaced by taking absolute values at both sides
Backtracking
$\textbf{p}_{k}^{\top}\nabla f_{k}\leq 0.$ (22)
A.0.2 Trust-region methods
Another class of optimization methods is the family of trust-region methods. The key distinction between trust-region and line-search methods is that the former defines a region (the trust-region) around the current point $\mathbf{x}_{k}$ and seeks a suitable point within this region that reduces the objective function. To achieve this, a trust-region algorithm first constructs a local quadratic approximation of the function, denoted as $m_{k}$, and then approximately solves the following subproblem:
subject to:
$\|p\|\leq\Delta_{k},$
$\textbf{x}_{k+1}=\textbf{x}_{k}+\textbf{p}_{k},$
that is, in a trust-region algorithm the direction and the step-length are calculated at the same time. The choice between backtracking line-search and the trust-region method depends on the problem characteristics and specific implementation details. Backtracking line-search is often preferable for simpler or computationally inexpensive problems due to its straightforward nature and adaptability. In contrast, trust-region methods are more suitable for complex or constrained optimization tasks where careful control over the step size is essential. A key distinction is that backtracking line-search directly adjusts the step size without solving quadratic subproblems, making it computationally faster in cases where objective function and gradient evaluations are inexpensive. Ultimately, the choice between these methods depends on the specific problem and computational trade-offs.
$\|\nabla f(\bm{x}_{k})\|\leq\epsilon,$ (23)
  1. 1.
  2. 2.
    If $\|\bm{x}_{k}-\bm{x}_{k-1}\|\leq\eta$, where $\eta$ is a small tolerance.
  3. 3.
$f(x,y)=x^{2}+y^{2}+xy,$
in Figures 26 and 27. We see that from iteration 1 to 6, the algorithm iteratively predicts the search direction. In this example, the trust-region radius is kept constant since $f(x)$ is convex. However, for PDE problems, we have implemented an adaptive trust-region size.
Appendix B Detailed Case Study: Burgers Equation in Single Precision
In this section, we provide additional details and case studies on the performance of quasi-Newton methods for solving the Burgers equation, with a particular focus on results obtained in single precision. We begin by examining the performance of SSBroyden, SSBFGS, and BFGS using Wolfe line-search.
Figure 28: PINNs with single precision for the Burgers equation. Results are shown with Wolfe line search. The plots show the evolution of the loss function (left) and the corresponding $l_{2}$ relative error (right) over training iterations for Case 1. After approximately 4,000 iterations, both the loss and error stabilize, indicating no significant further improvement.
Optimizer [# Iters.] Relative $l_{2}$ Error Training Time (s) Total Params
Adam [1000] + BFGS with Wolfe [10000] $1.04\times 10^{-4}$ 263 1,341
Adam [1000] + SSBroyden with Wolfe [10000] $3.60\times 10^{-5}$ 209 1,341
Adam [1000] + SSBFGS with Wolfe [10000] $4.04\times 10^{-5}$ 179 1,341
Table 17: PINNs with single precision for the Burgers equation. Relative $l_{2}$ error and training time for solving the Burgers equation using single precision with different optimizers. A PINN with four hidden layers, each containing 20 neurons, is trained for 1,000 iterations using the Adam optimizer, followed by 10,000 iterations using the specified quasi-Newton method.
Figure 29: Scenario 1: PINNs with single precision for the Burgers equation. Results shown with backtracking and trust-region strategies. The GPU is fully saturated with model parameters to minimize latency effects on compute time: The left panel illustrates a scenario where the iteration count of the first-order Adam algorithm dominates, combined with BFGS using the trust-region approach. This pairing exhibits degeneracy due to limited numerical precision during BFGS iterations. The right panel presents the convergence behavior of BFGS paired with trust-region and backtracking line search methods. In this setup, Adam is applied for 11 iterations as a warmup. The BFGS-based optimizer requires significantly more iterations than first-order optimization methods like Adam and Rectified Adam (RAdam) [79].
Adam [16001] + BFGS with trust-region [13] $2.91\times 10^{-3}$ 127 3501
Adam [11] + BFGS with backtracking [2113] $7.25\times 10^{-4}$ 61 3501
RAdam [11] + BFGS with backtracking [1008] $7.67\times 10^{-4}$ 62 3501
BFGS with backtracking [1677] $2.48\times 10^{-3}$ 45 3501
BFGS with trust-region [944] $1.39\times 10^{-3}$ 49 3501
Table 18: [Corresponding to Figure 29] PINNs with single precision for the Burgers equation: The relative $L_{2}$ error, training time, and number of training parameters are evaluated for solving the viscous Burgers equation with various optimizers and combinations of line search algorithms. The convergence criteria are based on absolute and relative tolerances in norm of the gradient of the loss function, which is set as $\text{ATOL},\text{RTOL}]=[10^{-7},10^{-7}]$. This experiments shows that Adam combined with BFGS trust-region provides better accuracy (hghlighted in bold fonts) .
Figure 30: Scenario 2: PINNs with single precision for the Burgers equation. Results are shown with backtracking and trust-region strategies. The GPU is fully saturated by collocation points and data points (initial and boundary conditions) to mitigate the impact of latency on computation time. The left panel depicts a scenario where the iteration count for the first-order Adam and RAdam optimizers is fixed at 1001, followed by the quasi-Newton BFGS optimizer using trust-region and backtracking line search algorithms. Similar to the results in Figure 29, this setup also exhibits degeneracy during BFGS iterations due to limited numerical precision. The right panel shows the convergence behavior of the standalone BFGS optimizer with trust-region and backtracking strategies, without any warmup from a first-order optimizer. In this configuration, the loss stagnates at $10^{-4}$ due to insufficient precision, particularly affecting the inversion of the Hessian matrix required for descent.
Adam [1001] + BFGS with trust-region [2142] $1.51\times 10^{-3}$ 30 1341
Adam [1001] + BFGS with backtracking [4544] $1.41\times 10^{-3}$ 43 1341
radam [1001] + BFGS with backtracking [4565] $1.43\times 10^{-3}$ 41 1341
BFGS with backtracking [2013] $1.65\times 10^{-3}$ 18 1341
BFGS with trust-region [2039] $\mathbf{1.1}\times\mathbf{10^{-3}}$ 24 1341
Table 19: [Corresponding to Figure 30] Performance metrics for the viscous Burgers equation using single-precision (32-bit) arithmetic for the cases shown in Figure 30: The relative $L_{2}$ error, training time, and number of training parameters are reported for solving the viscous Burgers equation with various optimizers and combinations of line search algorithms. The convergence criteria are defined by absolute and relative tolerances on the gradient norm of the loss function, set as $\text{ATOL},\text{RTOL}]=[10^{-7},10^{-7}]$. The results demonstrate that BFGS with the trust-region line search achieves the highest accuracy (highlighted in bold) among all cases presented in Figure 30.
Figure 28 illustrates the loss function and relative error over training iterations for SSBroyden, SSBFGS, and BFGS in single precision, corresponding to Case 1 in Table 1. The results are obtained using a PINN with four layers, each comprising 20 neurons, trained with the Adam optimizer followed by 10,000 iterations of the specified optimizer. Table 17 provides a summary of the corresponding relative error and training time. Notably, SSBFGS and SSBroyden achieve a relative error of $10^{-5}$, while BFGS converges to approximately $10^{-4}$. It is observed that the loss function and relative error stabilize after roughly 4,000 iterations, indicating no significant further improvement.
The performance of five different optimizer combinations is summarized in Table 18 for Scenario 1. The convergence history achieved using a neural network with 8 hidden layers, each containing 20 neurons, a $\tanh$ activation function, 200 randomly sampled data points for the initial and boundary conditions, and 10,000 collocation points to calculate the residual loss. The left panel of Figure 29 shows a scenario where the iteration count of the first-order Adam algorithm dominates, combined with BFGS using the trust-region line search. The key observation from this setup is that when Adam is run for an extended period, the BFGS optimizer provides no additional advantages due to the degeneration of the loss function caused by single-precision floating-point representation [Row 1 of Table 18]. However, when Adam or RAdam is used with a small number of iterations (11 in this case) as a warm-up followed by BFGS with both line search algorithms, the error improves by an order of magnitude [Rows 2, 3, and 4 of Table 18]. In contrast, starting directly with the BFGS optimizer results in errors of a similar magnitude to those in the first scenario (left panel of 29). All runs in Table 18 are terminated once the absolute and relative tolerances of $\text{ATOL},\text{RTOL}]=[10^{-7},10^{-7}]$ for the $L_{\infty}$ norm of the gradient of the loss function are met. Consequently, the number of iterations reported for the BFGS optimizers corresponds to the step where the specified tolerance is reached. The runtime for each case is also recorded in Table 18. For all runs, the runtime is less than one minute, except for the run where Adam is used for a larger number of iterations Figure (29).
In Scenario 2, where the GPU is fully utilized with collocation points and the PINN employs a smaller neural network, the convergence history for single precision arithmetic is shown in Figure 30. Convergence was achieved using a neural network with four hidden layers, each containing 20 neurons, a $\tanh$ activation function, 250 randomly sampled data points for initial and boundary conditions, and 50,000 collocation points. The performance metrics, including the relative $L_{2}$ error and training time for single and double precision, are summarized in Table 19.
Appendix C Lorenz system
@@
Appendix D Multi-dimensional Rosenbrock function